basicgamergirl:

PLS HELP / EXPLAIN !!!! The area of a right triangle is 180 m². The height of the right triangle is 40 m. What is the length of the hypotenuse of the right triangle? Enter your answer in the box. m

imqwerty:

\(\large Area= \frac{1}{2}(base)(height)\) you know the area and the height so you can find the base length using the formula and then use this to find the hypotenuse- \(\large (hypotenuse)^2=(base)^2+(height)^2\)